OpenStudy (anonymous):

Polynomial question Halp!!!!!!!!!! If f(x)= x^3 +6x^2-15x+c, find the two values of c and find the roots if two are identical

OpenStudy (anonymous):

Someone help plzzzzzzzzzzzzzzzz

ganeshie8 (ganeshie8):

say the roots are \(\alpha, \alpha ,\beta\)

ganeshie8 (ganeshie8):

\((x-\alpha)(x-\alpha)(x-\beta) = x^3 +6x^2-15x+c\) \((x-\alpha)^2(x-\beta) = x^3 +6x^2-15x+c\)

ganeshie8 (ganeshie8):

expand left side, and compare \(x^2, x\) coefficients, and constant terms you would get 3 equations in \(\alpha, \beta\) solve the equations and find \(\alpha, \beta\)